OpenStudy (logan13):

What is the length of stack M N with bar on top ? Round to the nearest tenth of a unit. units Line segment on a coordinate grid. Segment begins at point M, whose coordinates are two, negative two, and ends at point N, whose coordinates are eight, four.

OpenStudy (michele_laino):

hint: we have to do this computation: \[\begin{gathered} L = \sqrt {{{\left( {{x_N} - {x_M}} \right)}^2} + {{\left( {{y_N} - {y_M}} \right)}^2}} = \hfill \\ \hfill \\ = \sqrt {{{\left( {8 - 2} \right)}^2} + {{\left( {4 - \left( { - 2} \right)} \right)}^2}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (logan13):

uh.

OpenStudy (michele_laino):

such formula comes from the direct application of the Pythagorean Theorem

OpenStudy (michele_laino):

here are the next steps: \[\begin{gathered} \overline {MN} = \sqrt {{{\left( {{x_N} - {x_M}} \right)}^2} + {{\left( {{y_N} - {y_M}} \right)}^2}} = \hfill \\ \hfill \\ = \sqrt {{{\left( {8 - 2} \right)}^2} + {{\left( {4 - \left( { - 2} \right)} \right)}^2}} = \hfill \\ \hfill \\ = \sqrt {{6^2} + {{\left( {4 + 2} \right)}^2}} = \sqrt {{6^2} + {6^2}} = \hfill \\ \hfill \\ = \sqrt {36 + 36} = \sqrt {72} = ...? \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

so, what is the square root of 72?

OpenStudy (alivejeremy):

Can you write it for me?

OpenStudy (michele_laino):

yes! It is correct, nevertheless, the exercise asks for a decimal number rounded to the nearest tenth Now, we have: \[6\sqrt 2 = 8.48528...\] please round that number to the nearest tenth

OpenStudy (logan13):

8.4

OpenStudy (michele_laino):

better is \(8.5\) since the secon decimal figure is \(8\) which is greater than \(5\)

OpenStudy (michele_laino):

second*